OpenStudy (anonymous):

x-3/2x-4/3=-5/6(x+1/5)

OpenStudy (anonymous):

sorry, my mistake. multiply both sides by 30 \[30x-45x-40=-25(x+0,2)\] is what you'll get then\[10x-40=-5\] and in the end 10x=35 x=3,5

OpenStudy (anonymous):

I got \(x=\frac 7 2 \)